The ratio of the ages of Rajesh to that of Ramesh is 5 : 8. After 10 years the new ratio of their ages will be 7 : 10. What is the age of Ramesh at present?
Answer by lwsshak3(11628)

The ratio of the ages of Rajesh to that of Ramesh is 5 : 8. After 10 years the new ratio of their ages will be 7 : 10. What is the age of Ramesh at present?
**
let y=Rajesh's present age
let x=Ramesh's present age
y/x=5/8
y=5x/8
..
After 10 years:
y+10=Rajesh's age
x+10=Ramesh's age
(y+10)/(x+10)=7/10
7x+70=10y+100
7x+70=10(5x/8)+100
7x+50x/8=30
56x+50x=240
6x=240
x=40
Ramesh's present age=40